OpenStudy (anonymous):

find all complex numbers z satisfying z^3 = -27i in the form a + bi. Answers are posted below

OpenStudy (anonymous):

\[z = 3i\] \[z =\frac{ -3\sqrt{3} }{ 2 } - \frac{ 3i }{ 2}\] \[z = \frac{ 3\sqrt{3} }{ 2} - \frac{ 3i }{ 2 }\]